OpenStudy (anonymous):

Convert the decimal to a fraction.. Please show work... 0.24

OpenStudy (anonymous):

this is equal to 24/100. it needs reducing :)

OpenStudy (anonymous):

24/100=6/25

OpenStudy (anonymous):

0.24 = 24/100 = 6/25

OpenStudy (anonymous):

0.24 = 24/100. Then simplify by dividing the numerator (24) and denominator (100) of your fraction by the greatest common divisor: in this case 4. Thus, making (24/4)/(100/4) = 6/25.